worldwide emissions of carbon dioxide are estimated at 2x10^9 tons per year. what is this emission rate in pounds per hour
1
Expert's answer
2013-01-31T09:17:42-0500
1 Short Ton [US] = 2 000 Pounds

2x10^9 Ton = 4 * 10^13 Pounds

24 hours in a day times 365 days

24 x 365 = 8760 hours

Emission rate in pounds per hour = 4 * 10^13 / 8760 = 4.57 * 10^9 pounds per hour